OpenStudy (anonymous):

5. A private school opened in 2005 with an initial enrollment of 85 students. The enrollment has increased by an average of 18 students each year since the school opened. A nearby public school had an enrollment of 95 in 2005, and its enrollment has increased by an average of 15 students per year. 1. Let t = time in years, with t = 0 representing the year 2005. Let f(t) = the number of students enrolled at the private school and g(t) = the number of students enrolled at the public school. Create the two functions to represent the situation. f(t)+ 18 g(t) + 15 2. Graph the system of equation.

OpenStudy (anonymous):

.5. A private school opened in 2005 with an initial enrollment of 85 students. The enrollment has increased by an average of 18 students each year since the school opened. A nearby public school had an enrollment of 95 in 2005, and its enrollment has increased by an average of 15 students per year. 1. Let t = time in years, with t = 0 representing the year 2005. Let f(t) = the number of students enrolled at the private school and g(t) = the number of students enrolled at the public school. Create the two functions to represent the situation. f(t)+ 18 g(t) + 15 2. Graph the system of equations to determine the point of intersection. 85(1) + 18 85( 95(1) +15 3. Interpret the point of intersection in the context of the problem.

OpenStudy (anonymous):

i need help for 3..

OpenStudy (anonymous):

and 2

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

OK. We did this last night. private school: f(t) = 85 +18t public school: g(t) = 95 + 15t For 2, you can just rewrite them as y = 85 + 18x y = 95 + 15x And graph that

OpenStudy (bibby):

for 3, to find the intersection of any 2 graphs, you set them equal to each other. Solve for x: 95 + 15x = 85 + 18x

OpenStudy (anonymous):

ok so thats..... it will take a while.

OpenStudy (bibby):

lol what are you having problems with?

OpenStudy (anonymous):

well trying to find the right number that they both meet.

OpenStudy (bibby):

you don't have to guess. Just solve for x 95 + 15x = 85 + 18x Start by getting the x's onto one side

OpenStudy (anonymous):

So i subtract 15x from 18x and get 3x ?

OpenStudy (bibby):

Exactly. But you have to subtract 15x from the left too. 95 + 15x = 85 + 18x -15x -15x --------------------- 95 = 85 + 3x

OpenStudy (anonymous):

then i subtract 85 from 95 and get 10 x = 10?

OpenStudy (bibby):

what about the 3x?

OpenStudy (anonymous):

ohh isnt it 7= x or 13 = x

OpenStudy (bibby):

95 = 85 + 3x -85 -85 ------ 10 = 3x

OpenStudy (anonymous):

so we have to subtract 3 from each?

OpenStudy (bibby):

it's 3x not 3. 3 times x. Not 3+x 10=3x -> divide by 3 on both sides

OpenStudy (anonymous):

ohhh yeaa! 10/3 = 3.3333333.......

OpenStudy (bibby):

you can just put it in as 10/3 I think.

OpenStudy (anonymous):

ok thank you .. i have another that i need help on.

OpenStudy (anonymous):

but 10/3 as a year??

OpenStudy (bibby):

I guess you're right. Either way, they want the point, that's not the final answer. We have to plug in x=10/3 into one of the equation

OpenStudy (anonymous):

ok 85 + 18 x 10/3

OpenStudy (anonymous):